# A scattering perspective on gravitational lensing

### v1

@@ Abstract
Introduction. The advent of gravitational-wave (GW) astronomy has enabled direct studies of compact objects and their populations across cosmic time. Since the first detection of a binary black hole merger in 2015 LIGOScientific:2016aoc, the LIGO–Virgo–KAGRA Collaboration has progressively expanded its sensitivity and frequency coverage LIGOScientific:2018mvr; LIGOScientific:2020ibl; KAGRA:2021vkt, recently releasing its fourth observing run catalog LIGOScientific:2025hdt; LIGOScientific:2025slb. Together with the three preceding observing runs, the full catalog now includes hundreds of confirmed compact-binary coalescences, spanning a wide range of source masses, mass ratios, and redshifts. As the number of detections grows, especially with the advent of next generation detectors such as the Einstein Telescope, Cosmic Explorer and LISA ET:2025xjr; Branchesi:2023mws; LISA:2024hlh, the probability that some of these signals have been affected by gravitational lensing also increases, making wave propagation effects through intervening structures an important ingredient in the interpretation of current and future GW data Hannuksela:2019kle; Liu:2020par; LIGOScientific:2021izm; Jung:2017flg; Wempe:2022zlk; LIGOScientific:2023bwz; Oguri:2018muv; Yin:2023kzr; Sereno:2010dr; Lai:2018rto; Urrutia:2021qak; Baker:2016reh. Intriguingly, the recently reported event GW231123 has been highlighted as a potential candidate for providing observational evidence of GW lensing LIGOScientific:2025rsn.
@@
Depending on the ratio between the GW wavelength and the characteristic scale of the lens potential, gravitational lensing operates in qualitatively different regimes. In the geometric-optics limit, where the wavelength is negligible compared to the lens scale, the signal is described as a sum of distinct images associated with stationary paths of the Fermat potential, each characterized by a magnification and a time delay Schneider:1992bmb; Meneghetti:2020yif. In contrast, when the wavelength becomes comparable to the Schwarzschild radius of the lens, wave effects become dominant and diffraction and interference imprint frequency-dependent modulations on the observed waveform Ohanian:1974ys; Bliokh; Bontz:1981rvr; Deguchi:1986zz; Takahashi:2003ix; Schneider:1992bmb; Nakamura:1999uwi (for a recent review on gravitational lensing in the context of GWs, see Ref. Grespan:2023cpa).
@@
A general treatment of GW lensing in the wave regime is provided by the diffraction integral, which captures the coherent superposition of all propagation paths through the lens plane Takahashi:2003ix; Oguri:2020ldf; Nakamura:1999uwi; Takahashi:2004mc; Takahashi:2005sxa; Takahashi:2005ug. This formulation contains both the geometric and diffractive limit as special cases: in the high-frequency limit, the integral reduces to a discrete sum over stationary paths corresponding to the standard image positions, while in the low-frequency limit it describes a single, smooth amplification factor accounting for interference and phase modulation Jow:2022pux. The diffraction integral thus provides a unified mathematical description of GW lensing that naturally bridges the wave and geometric regimes Takahashi:2003ix; Oguri:2020ldf; Nakamura:1999uwi; Takahashi:2004mc; Takahashi:2005sxa; Takahashi:2005ug.
@@
Given the upcoming observational prospects, the literature focusing on wave-optics effects has been growing. Several groups have developed numerical codes to compute the diffraction integral efficiently yeung2024wolensing; Villarrubia-Rojo:2024xcj; Feldbrugge:2019fjs. These codes have been used to study the properties of different lens models Tambalo:2022wlm; Caliskan:2022hbu, including rotating lenses Bonga:2024orc; Kubota:2023dlz; Kubota:2023dlz; Li:2022izh, lenses in binary systems Feldbrugge:2020ycp; Mehrabi:2012dy or the intriguing possibility that the lens is given by a different dark matter model Singh:2025uvp. Wave-optics effects also play an important role in placing constraints on the abundance of primordial black holes Sugiyama:2019dgt. With a particular focus on the space-based interferometer LISA, probabilities of observing such effects and parameter estimations for the lens properties can be found in Caliskan:2023zqm; Caliskan:2022hbu.
@@
In this paper, we address this issue by connecting the diffraction integral description of lensing to the more general framework of scattering theory. Although the analogy between gravitational lensing and wave scattering has long been recognized, the formal connection between the diffraction integral formalism and the general theory of wave propagation—such as that provided by scattering theory Newton:1982qc—has not been systematically explored. Yet, the structure of a lensing event closely mirrors that of a scattering process, in which an incoming wave interacts with an external potential and emerges with a modified amplitude and phase Futterman:1988ni; Vishveshwara:1970zz; Press:1971wr; Macquart:2004sh. This perspective suggests that gravitational lensing can be understood as the scattering of gravitational waves by the gravitational potential of the lens, offering a conceptually transparent and physically unified framework with well-defined expansion parameters governing the validity of the results Virbhadra:1999nm; Bozza:2001xd; Bozza:2002af; Pijnenburg:2024btj; Chan:2025wgz; Yin:2023kzr; Motohashi:2021zyv. More specifically, gravitational lensing corresponds to the gravitational Compton scattering process, which has largely been studied in the on-shell program for gravitational waves from binary systems Bjerrum-Bohr:2014zsa; Bjerrum-Bohr:2016hpa; Bai:2016ivl; Chi:2019owc; AccettulliHuber:2020oou; Bastianelli:2021nbs; Chen:2022clh; Comberiati:2024uuc; Cangemi:2023bpe; Ivanov:2024sds; Correia:2024jgr; Caron-Huot:2025tlq.
@@
In the standard lensing framework, the lens is modeled as a Newtonian, static potential (later, we will relax this assumption to incorporate relativistic corrections). Despite the spin-2 nature of GWs, we will consider the simple case of scalar fields, hence phenomena such as the mixing of different polarization modes Andersson:2020gsj; Oancea:2022szu will be ignored. While these corrections might be relevant in the Born regime Garoffolo:2022usx, they are suppressed in the eikonal limit, which we adopt in many parts of this work (see the Beyond eikonal corrections Section for additional details). Therefore, the propagation equation reads Grespan:2023cpa
@@
up to first order in the gravitational potential of the lens, $U_{\rm L}$, and where we have performed a Fourier transformation to the frequency domain, with mode frequency $\omega$. For distant GW sources, the emission can be approximated as point-like, so that the unlensed signal is typically modeled as a spherical wave of the form $\phi^{0}_{\omega,{\rm L}}({\bm{r}})=e^{i\omega|{\bm{r}}-{\bm{r}}_{\rm s}|}/|{\bm{r}}-{\bm{r}}_{\rm s}|$. By introducing an amplification factor $F({\bm{r}})$ for the full lensed wave,
@@
written in spherical coordinates $\{\tilde{\theta},\tilde{\varphi}\}$ centered on the source, with $\tilde{r}=|{\bm{r}}-{\bm{r}}_{\rm s}|$ and $\nabla^{2}_{\tilde{\bm{\theta}}}$ the Laplacian on the sphere. As shown in Nakamura:1999uwi (see also Braga:2024pik; Feldbrugge:2019fjs for alternative derivations), one solution of Eq. (3) can be found under the following set of assumptions, namely
@@
Eq. (4) expresses the paraxial (or small-angle) approximation, in which the wave propagates predominantly along a single direction (taken here as the $z$-axis), such that it makes only small angles with that axis, $\tilde{\bm{\theta}}\ll 1$. [^2] When expanding the spherical Laplacian in the paraxial approximation, spherical coordinates reduce to cylindrical ones, so that centering the $z$-axis on the source or on the lens is equivalent. This also allows to consider angles centered on the lens, $\{\theta,\varphi\}$. On the other hand, Eq. (5) is typically identified as the eikonal approximation in the lensing literature, despite its practical interpretation often being ambiguous. As we show in this work, this assumption is equivalent to the conventional eikonal approximation used in scattering theory, corresponding to the regime of large impact parameters compared to the GW wavelength. These assumptions allow to convert the wave equation into a Schrödinger-like form, which can be solved in terms of a path integral according to the expression Nakamura:1999uwi
@@
where the label $I$ indicates the (potentially multiple) “images” associated to the stationary points of the phase. Then, the diffraction integral simplifies to the form Nakamura:1999uwi; Grespan:2023cpa
@@
For the lensing framework at hand, the scattering of a gravitational wave by a localized lens can be modeled as the Compton scattering of a graviton or, in this case, a massless scalar with a point particle, described in terms of an effective theory with field Cheung:2018wkq; Bern:2019crd; Bern:2019nnu or worldline Goldberger:2004jt; Mogull:2020sak degrees of freedom. Following this approach, the corresponding quantum field theory scattering amplitude $\mathcal{M}$ can be related in the classical limit to the scattering amplitude $f$ as [^3] We adopt the normalization of Correia:2024jgr, which ensures consistency between the Lippmann–Schwinger and the Schrödinger-like equation used in the lensing case.
@@
where $M$ is the lens mass, related to its gravitational potential $U_{\rm L}$ as discussed below, while $b=|{\bm{b}}|$. Following these assumptions, the classical amplitude will be expressed as a series expansion in $GM/b$, encoding the classical non-linearities. The scattering amplitude and lens potential are then related by the Lippmann-Schwinger equation Iwasaki:1971vb; Cristofoli:2019neg which, in momentum space, reads
@@
Here, the incoming graviton and the lens have four-momenta $p_{1}^{\mu}=(\omega,{\bm{p}})$ and $p_{2}^{\mu}=(\sqrt{\omega^{2}+M^{2}},-{\bm{p}})$, respectively, such that $\sqrt{s}\simeq M$. We have also defined $\int_{\bm{k}}\equiv\int{\rm d}^{3}{\bm{k}}/(2\pi)^{3}$. Notice that the momentum-space potential $\mathcal{V}$ entering in the Lippmann-Schwinger equation is defined through the Fourier transform Correia:2024jgr
@@
in terms of its coordinate-space counterpart $V(\mathbf{r})$, which in turn is related to that defined in Eq. (1) by $V({\bm{r}})=4\omega^{2}U_{\rm L}({\bm{r}})$. Given the form of $V$, one can first determine $\mathcal{V}(\bm{p},\bm{p}+\bm{q})$, and then solve the Lippmann-Schwinger equation (A scattering perspective on gravitational lensing) perturbatively.
@@
On the lensing side, the Born approximation is equivalent to expanding the exponential in Eq. (7) for small potentials $\psi$, yielding Nakamura:1999uwi
@@
where we have expanded the time delay $t_{d}$, used Eq. (20) to neglect the term $\propto{\bm{b}}^{2}$, and defined the Fresnel scale [^4] Notice that the usual Fresnel scale is $r_{\rm F}^{\prime}=\sqrt{z_{\rm sl}z_{\rm lo}/(z_{\rm so}\omega)}$, while here we have defined $r_{\rm F}$ as its projection onto the observer plane., $r_{\rm F}=\sqrt{z_{\rm so}z_{\rm lo}/(z_{\rm sl}\omega)}$. Using Eq. (14), the plane wave in the expression (21) can also be written as $e^{-\frac{i\omega{\bm{b}}\cdot{\bm{b}}_{\rm o}}{z_{\rm lo}}}=e^{-i{\bm{b}}\cdot{\bm{p}}^{\prime}}=e^{-i{\bm{b}}\cdot{\bm{q}}}$, if one promotes $\bm{b}$} to a 3D vector with null $z$-component, since $\bm{p}$} and $\bm{b}$} are orthogonal to each other. See Yarimoto:2024uew for a systematic treatment beyond the leading-order Born approximation.
@@
First, we fix the relative normalization $\cal N$} by matching the free waves close to the scattering surface (i.e. far away from the source, where the spherical wave has flat wavefronts). This yields Pijnenburg:2024btj
@@
which is the expected result of the scattering amplitude in the Born approximation, obtained as the lowest order solution of the Lippmann-Schwinger equation (A scattering perspective on gravitational lensing).
@@
where we have introduced the dimensionless frequency $\nu=4GM\omega$, the Einstein radius $r^{2}_{\rm E}=4GMz_{\rm sl}z_{\rm lo}/z_{\rm so}$, and the impact parameter on the lens plane ${\bm{b}}_{\rm l}=(z_{\rm sl}/z_{\rm so}){\bm{b}}_{\rm o}$. Our result matches the one in Takahashi:2005sxa in the limit $(r_{\rm E}/|{\bm{b}}_{\rm l}|)^{2}\ll\nu$ or, equivalently, $|{\bm{q}}||{\bm{b}}_{\rm l}|\gg 1$.
Partial waves and the Fresnel factor. In the next two sections we will demonstrate that a similar relation between the scattering amplitude and the amplification factor can also be derived in the context of a partial wave analysis. Indeed, it is known that, for central potentials, Eq. (1) can be solved by performing a decomposition over spherical harmonics. This is the standard lore in scattering theory, leading to the usual form of $f$ given in terms of the phase shifts $\delta_{\ell}$ Futterman:1988ni; Andersson:1995vi; Glampedakis:2001cx
@@
where $r$ and $r_{\rm s}$ denote the relative distances between the lens and the observer or source, respectively, with $r<r_{\rm s}$, [^5] Notice that choosing $r>r_{s}$ would demand swapping the arguments of the spherical Bessel and Hankel functions, yielding an equivalent result. while the factor $(-1)^{\ell}$ arises from setting the source on the z-z axis. Comparing Eqs. (30) and (31), one can see that the only difference arises from the additional Hankel function $h^{(1)}_{\ell}(\omega r_{\rm s})$, if the incoming wave is spherical. Then, one can separate $j_{\ell}(\omega r)$ into its two spherical Hankel components, and collect the outgoing components of the full wave at large distances, to obtain Nambu:2015aea; Motohashi:2021zyv [^6] The phase shifts $\delta_{\ell}$ are defined in terms of the asymptotic radial behavior of the scalar field according to Futterman:1988ni $\phi_{\omega}({\bm{r}})\propto r^{-1}e^{-i\omega r}+(-1)^{\ell+1}e^{2i\delta_{\ell}}\,r^{-1}e^{i\omega r}\,,\qquad r\to\infty\,$, (32) when neglecting absorption processes.
@@
Extracting the scattering amplitude $f$ from this expression requires taking the large distance limit, by approximating the spherical Hankel functions as $h^{(1)}_{\ell}(x)\xrightarrow[x\to\infty]{}(-i)^{\ell+1}e^{ix}/x$. At first order, one finds that the difference between an incoming plane or spherical wave thus reduces to an overall factor $e^{i\omega r_{\rm s}}/\omega r_{\rm s}$ that can be reabsorbed into the normalization of the total wave. However, since lensing results are typically obtained in the finite-distance limit, we include one additional order in the expansion of the spherical Hankel functions by considering Turyshev:2021rkj; Turyshev:2018gjj
@@
In other words, through the inclusion of this correction, the scattering amplitude retains a nontrivial radial dependence, in analogy to the diffraction integral, allowing for a matching prescription. Furthermore, let us stress that, as discussed in Ref. Nambu:2015aea, one could obtain the quadratic term of Eq. (34) by considering a WKB solution to the wave equation. When following this approach, it is possible to improve the small $\ell$ behavior through the replacement $\ell(\ell+1)\to(\ell+1/2)^{2}$, as shown by Langer Langer:1937qr, getting the final expression
@@
This is nothing but the standard result for the eikonal scattering between a massless particle [^7] In the eikonal limit, the scalar field scattering amplitude reproduces the gravitational one, while subleading terms encode graviton helicity-flip effects (see comments in the Conclusions). and a massive (lens) field, obtained by resumming all the ladder and cross-ladder diagrams in its Feynman series expansion Abarbanel:1969ek; Levy:1969cr; tHooft:1987vrq, with the addition of the finite distance Fresnel term. The phase shift in the eikonal limit is then given by Abarbanel:1969ek; Levy:1969cr; tHooft:1987vrq
@@
where $\mathcal{M}|_{G}({\bm{q}},s)$ describes the tree-level $t$-channel diagram. In the second line, we have used that $2(s-M^{2})\simeq 4M\omega$, the leading order Lippmann-Schwinger equation (A scattering perspective on gravitational lensing) to write the scattering amplitude in terms of the lens potential, and the Fourier transform conventions in Eq. (19). Thus, comparing to Eq. (9), one derives
@@
reproducing the known result that the two dimensional potential $\psi$ is equal to the gravitational time delay, $\partial_{\omega}\delta$ Schneider:1992bmb. In particular, for the potential of a Newtonian point-like lens, the eikonal phase shift takes the simple form
@@
which reproduces the matching relation (24) obtained in the Born approximation. This matching clarifies the crucial assumption of Eq. (5), $|\partial_{\tilde{r}}^{2}F|\ll|\omega\partial_{\tilde{r}}F|$, which amounts to adopting the eikonal approximation $\ell\approx\omega b\gg 1$ (together with small deflection angles (4)) as essential in drawing an analogy between the lensing diffraction integral and scattering frameworks, in addition to the condition of classicality listed in Eq. (17).
Finally, we note that the unitarity properties of the scattering amplitude allow one to derive optical-theorem–type statements for the diffraction integral (for work along these lines, but focusing on the consequences of causality, see Tanaka:2023mvy; Tanaka:2025ntr; Suyama:2025gbh). Since, by construction, the eikonal amplitude satisfies the optical theorem, the amplification factor given by the diffraction integral satisfies [^8] Here we are assuming conservative potentials, but this can be easily generalized to include inelastic effects.

### v2

# Scattering perspective on gravitational lensing

@@ Abstract
Introduction. The advent of gravitational-wave (GW) astronomy has enabled direct studies of compact objects and their populations across cosmic time. Since the first detection of a binary black hole merger in 2015 Abbott and others (2016), the LIGO-Virgo-KAGRA Collaboration has progressively expanded its sensitivity and frequency coverage Abbott and others (2019, 2021a, 2023), recently releasing its fourth observing run catalog Abac and others (2025b, c). Together with the three preceding observing runs, the full catalog now includes hundreds of confirmed compact-binary coalescences, spanning a wide range of source masses, mass ratios, and redshifts. As the number of detections grows, especially with the advent of next generation detectors such as the Einstein Telescope, Cosmic Explorer and LISA Abac and others (2025d); Branchesi and others (2023); Colpi and others (2024), the probability that some of these signals have been affected by gravitational lensing also increases, making wave propagation effects through intervening structures an important ingredient in the interpretation of current and future GW data Hannuksela et al. (2019); Liu et al. (2021); Abbott and others (2021b); Jung and Shin (2019); Wempe et al. (2024); Abbott and others (2024); Oguri (2018); Yin and He (2024); Sereno et al. (2010); Lai et al. (2018); Urrutia and Vaskonen (2021); Baker and Trodden (2017). Intriguingly, the recently reported event GW231123 has been highlighted as a potential candidate for providing observational evidence of GW lensing Abac and others (2025a).
@@
Depending on the ratio between the GW wavelength and the characteristic scale of the lens potential, gravitational lensing operates in qualitatively different regimes. In the geometric-optics limit, where the wavelength is negligible compared to the lens scale, the signal is described as a sum of distinct images associated with stationary paths of the Fermat potential, each characterized by a magnification and a time delay Schneider et al. (1992); Meneghetti and others (2020). In contrast, when the wavelength becomes comparable to the Schwarzschild radius of the lens, wave effects become dominant and diffraction and interference imprint frequency-dependent modulations on the observed waveform Ohanian (1974); Bliokh and Minakov (1975); Bontz and Haugan (1981); Deguchi and Watson (1986); Takahashi and Nakamura (2003); Schneider et al. (1992); Nakamura and Deguchi (1999) (for a recent review on gravitational lensing in the context of GWs, see Ref. Grespan and Biesiada (2023)).
@@
A general treatment of GW lensing in the wave regime is provided by the diffraction integral, which captures the coherent superposition of all propagation paths through the lens plane Takahashi and Nakamura (2003); Oguri and Takahashi (2020); Nakamura and Deguchi (1999); Takahashi (2004); Takahashi et al. (2005); Takahashi (2006). This formulation contains both the geometric and diffractive limits as special cases: in the high-frequency limit, the integral reduces to a discrete sum over stationary paths corresponding to the standard image positions, while in the low-frequency limit it describes a single, smooth amplification factor accounting for interference and phase modulation Jow et al. (2023). The diffraction integral thus provides a unified mathematical description of GW lensing that naturally bridges the wave and geometric regimes Takahashi and Nakamura (2003); Oguri and Takahashi (2020); Nakamura and Deguchi (1999); Takahashi (2004); Takahashi et al. (2005); Takahashi (2006).
@@
Given the upcoming observational prospects, the literature focusing on wave-optics effects has been growing. Several groups have developed numerical codes to compute the diffraction integral efficiently Yeung et al. (2024); Villarrubia-Rojo et al. (2025); Feldbrugge et al. (2023); Diego et al. (2019). These codes have been used to study the properties of different lens models Tambalo et al. (2023); Çalışkan et al. (2023b), including rotating lenses Bonga et al. (2025); Kubota et al. (2024b, b); Li et al. (2022), lenses in binary systems Feldbrugge and Turok (2020); Mehrabi and Rahvar (2013) or the intriguing possibility that the lens is given by a different dark matter model Singh et al. (2025). Wave-optics effects also play an important role in placing constraints on the abundance of primordial black holes Sugiyama et al. (2020); Diego (2020). With a particular focus on the space-based interferometer LISA, probabilities of observing such effects and parameter estimations for the lens properties can be found in Çalışkan et al. (2023a, b).
@@
In this paper, we address this issue by connecting the diffraction integral description of lensing to the more general framework of scattering theory. Although the analogy between gravitational lensing and wave scattering has long been recognized, the formal connection between the diffraction integral formalism and the general theory of wave propagation—such as that provided by scattering theory Newton (1982)—has not been systematically explored. Yet, the structure of a lensing event closely mirrors that of a scattering process, in which an incoming wave interacts with an external potential and emerges with a modified amplitude and phase Futterman et al. (2012); Vishveshwara (1970); Press (1971); Macquart (2004). This perspective suggests that gravitational lensing can be understood as the scattering of gravitational waves by the gravitational potential of the lens, offering a conceptually transparent and physically unified framework with well-defined expansion parameters governing the validity of the results Virbhadra and Ellis (2000); Bozza et al. (2001); Bozza (2003); Pijnenburg et al. (2024); Chan et al. (2025); Yin and He (2024); Motohashi and Noda (2021). More specifically, gravitational lensing corresponds to the gravitational Compton scattering process, which has largely been studied in the on-shell program for gravitational waves from binary systems Bjerrum-Bohr et al. (2015, 2016); Bai and Huang (2017); Chi (2019); Accettulli Huber et al. (2020); Bastianelli et al. (2022); Chen et al. (2022a); Comberiati and de la Cruz (2025); Cangemi et al. (2024); Ivanov et al. (2024); Correia and Isabella (2025); Caron-Huot et al. (2025); Bautista and Siemonsen (2022); Bautista et al. (2023a, b); Bjerrum-Bohr et al. (2025); Akpinar (2025); Di Vecchia et al. (2024); Koemans Collado et al. (2019); Aoude et al. (2022a); Chen et al. (2022b); Chiodaroli et al. (2022); Aoude et al. (2022b); Bjerrum-Bohr et al. (2023, 2024); Vazquez-Holm and Luna (2025).
@@
In the standard lensing framework, the lens is modeled as a Newtonian, static potential (later, we will relax this assumption to incorporate relativistic corrections). Despite the spin-2 nature of GWs, we will consider the simple case of scalar fields, hence phenomena such as the mixing of different polarization modes Andersson et al. (2021); Oancea et al. (2024) will be ignored. While these corrections might be relevant in the Born regime Garoffolo (2024), they are suppressed in the eikonal limit, which we adopt in many parts of this work (see the Beyond eikonal corrections Section for additional details). Therefore, the propagation equation reads Grespan and Biesiada (2023)
@@
up to first order in the gravitational potential of the lens, $U_{\rm L}$, and where we have performed a Fourier transformation to the frequency domain, with mode frequency $\omega$. For distant GW sources, the emission can be approximated as pointlike, so that the unlensed signal is typically modeled as a spherical wave of the form $\phi^{0}_{\omega,{\rm L}}({\bm{r}})=e^{i\omega|{\bm{r}}-{\bm{r}}_{\rm s}|}/|{\bm{r}}-{\bm{r}}_{\rm s}|$. By introducing an amplification factor $F({\bm{r}})$ for the full lensed wave,
@@
written in spherical coordinates $\{\tilde{\theta},\tilde{\varphi}\}$ centered on the source, with $\tilde{r}=|{\bm{r}}-{\bm{r}}_{\rm s}|$ and $\nabla^{2}_{\tilde{\bm{\theta}}}$ the Laplacian on the sphere. As shown in Nakamura and Deguchi (1999) (see also Braga et al. (2024); Feldbrugge et al. (2023) for alternative derivations), one solution of Eq. (3) can be found under the following set of assumptions, namely
@@
Eq. (4) expresses the paraxial (or small-angle) approximation, in which the wave propagates predominantly along a single direction (taken here as the $z$-axis), such that it makes only small angles with that axis, $\tilde{\bm{\theta}}\ll 1$. [^2] When expanding the spherical Laplacian in the paraxial approximation, spherical coordinates reduce to cylindrical ones, so that centering the $z$-axis on the source or on the lens is equivalent. This also allows one to consider angles centered on the lens, $\{\theta,\varphi\}$. On the other hand, Eq. (5) is typically identified as the eikonal approximation in the lensing literature, despite its practical interpretation often being ambiguous. As we show in this work, this assumption is equivalent to the conventional eikonal approximation used in scattering theory, corresponding to the regime of large impact parameters compared to the GW wavelength. These assumptions allow one to convert the wave equation into a Schrödinger-like form, which can be solved in terms of a path integral according to the expression Nakamura and Deguchi (1999)
@@
where the label $I$ indicates the (potentially multiple) “images” associated to the stationary points of the phase. Then, the diffraction integral simplifies to the form Nakamura and Deguchi (1999); Grespan and Biesiada (2023)
@@
For the lensing framework at hand, the scattering of a gravitational wave by a localized lens can be modeled as the Compton scattering of a graviton or, in this case, a massless scalar with a point particle, described in terms of an effective theory with field Cheung et al. (2018); Bern et al. (2019a, b) or worldline Goldberger and Rothstein (2006); Mogull et al. (2021) degrees of freedom. Following this approach, the corresponding quantum field theory scattering amplitude $\mathcal{M}$ can be related in the classical limit to the scattering amplitude $f$ as [^3] We adopt the normalization of Correia and Isabella (2025), which ensures consistency between the Lippmann-Schwinger and the Schrödinger-like equation used in the lensing case.
@@
where $M$ is the lens mass, related to its gravitational potential $U_{\rm L}$ as discussed below, while $b=|{\bm{b}}|$. Following these assumptions, the classical amplitude will be expressed as a series expansion in $GM/b$, encoding the classical non-linearities. The scattering amplitude and lens potential are then related by the Lippmann-Schwinger equation Iwasaki (1971); Cristofoli et al. (2019) which, in momentum space, reads
@@
Here, the incoming graviton and the lens have four-momenta $p_{1}^{\mu}=(\omega,{\bm{p}})$ and $p_{2}^{\mu}=(\sqrt{\omega^{2}+M^{2}},-{\bm{p}})$, respectively, such that $\sqrt{s}\simeq M$. We have also defined $\int_{\bm{k}}\equiv\int{\rm d}^{3}{\bm{k}}/(2\pi)^{3}$. Notice that the momentum-space potential $\mathcal{V}$ entering in the Lippmann-Schwinger equation is defined through the Fourier transform Correia and Isabella (2025)
@@
in terms of its coordinate-space counterpart $V(\mathbf{r})$, which in turn is related to that defined in Eq. (1) by $V({\bm{r}})=4\omega^{2}U_{\rm L}({\bm{r}})$. Given the form of $V$, one can first determine $\mathcal{V}(\bm{p},\bm{p}+\bm{q})$, and then solve the Lippmann-Schwinger equation (Scattering perspective on gravitational lensing) perturbatively.
@@
On the lensing side, the Born approximation is equivalent to expanding the exponential in Eq. (7) for small potentials $\psi$, yielding Nakamura and Deguchi (1999)
@@
where we have expanded the time delay $t_{d}$, used Eq. (20) to neglect the term $\propto{\bm{b}}^{2}$, and defined the Fresnel scale [^4] Notice that the usual Fresnel scale is $r_{\rm F}^{\prime}=\sqrt{z_{\rm sl}z_{\rm lo}/(z_{\rm so}\omega)}$, while here we have defined $r_{\rm F}$ as its projection onto the observer plane., $r_{\rm F}=\sqrt{z_{\rm so}z_{\rm lo}/(z_{\rm sl}\omega)}$. Using Eq. (14), the plane wave in the expression (21) can also be written as $e^{-\frac{i\omega{\bm{b}}\cdot{\bm{b}}_{\rm o}}{z_{\rm lo}}}=e^{-i{\bm{b}}\cdot{\bm{p}}^{\prime}}=e^{-i{\bm{b}}\cdot{\bm{q}}}$, if one promotes $\bm{b}$} to a 3D vector with null $z$-component, since $\bm{p}$} and $\bm{b}$} are orthogonal to each other. See Yarimoto and Oguri (2025) for a systematic treatment beyond the leading-order Born approximation.
@@
First, we fix the relative normalization $\cal N$} by matching the free waves close to the scattering surface (i.e. far away from the source, where the spherical wave has flat wavefronts). This yields Pijnenburg et al. (2024)
@@
which is the expected result of the scattering amplitude in the Born approximation, obtained as the lowest order solution of the Lippmann-Schwinger equation (Scattering perspective on gravitational lensing).
@@
where we have introduced the dimensionless frequency $\nu=4GM\omega$, the Einstein radius $r^{2}_{\rm E}=4GMz_{\rm sl}z_{\rm lo}/z_{\rm so}$, and the impact parameter on the lens plane ${\bm{b}}_{\rm l}=(z_{\rm sl}/z_{\rm so}){\bm{b}}_{\rm o}$. Our result matches the one in Takahashi et al. (2005) in the limit $(r_{\rm E}/|{\bm{b}}_{\rm l}|)^{2}\ll\nu$ or, equivalently, $|{\bm{q}}||{\bm{b}}_{\rm l}|\gg 1$.
Partial waves and the Fresnel factor. In the next two sections we will demonstrate that a similar relation between the scattering amplitude and the amplification factor can also be derived in the context of a partial wave analysis. Indeed, it is known that, for central potentials, Eq. (1) can be solved by performing a decomposition over spherical harmonics. This is the standard lore in scattering theory, leading to the usual form of $f$ given in terms of the phase shifts $\delta_{\ell}$ Futterman et al. (2012); Andersson (1995); Glampedakis and Andersson (2001)
@@
where $r$ and $r_{\rm s}$ denote the relative distances between the lens and the observer or source, respectively, with $r<r_{\rm s}$, [^5] Notice that choosing $r>r_{s}$ would demand swapping the arguments of the spherical Bessel and Hankel functions, yielding an equivalent result. while the factor $(-1)^{\ell}$ arises from setting the source on the z-z axis. Comparing Eqs. (30) and (31), one can see that the only difference arises from the additional Hankel function $h^{(1)}_{\ell}(\omega r_{\rm s})$, if the incoming wave is spherical. Then, one can separate $j_{\ell}(\omega r)$ into its two spherical Hankel components, and collect the outgoing components of the full wave at large distances, to obtain Nambu and Noda (2016); Motohashi and Noda (2021) [^6] The phase shifts $\delta_{\ell}$ are defined in terms of the asymptotic radial behavior of the scalar field according to Futterman et al. (2012) $\phi_{\omega}({\bm{r}})\propto r^{-1}e^{-i\omega r}+(-1)^{\ell+1}e^{2i\delta_{\ell}}\,r^{-1}e^{i\omega r}\,,\qquad r\to\infty\,$, (32) when neglecting absorption processes.
@@
Extracting the scattering amplitude $f$ from this expression requires taking the large distance limit, by approximating the spherical Hankel functions as $h^{(1)}_{\ell}(x)\xrightarrow[x\to\infty]{}(-i)^{\ell+1}e^{ix}/x$. At first order, one finds that the difference between an incoming plane or spherical wave thus reduces to an overall factor $e^{i\omega r_{\rm s}}/\omega r_{\rm s}$ that can be reabsorbed into the normalization of the total wave. However, since lensing results are typically obtained in the finite-distance limit, we include one additional order in the expansion of the spherical Hankel functions by considering Turyshev and Toth (2021, 2019)
@@
In other words, through the inclusion of this correction, the scattering amplitude retains a nontrivial radial dependence, in analogy to the diffraction integral, allowing for a matching prescription. Furthermore, let us stress that, as discussed in Ref. Nambu and Noda (2016), one could obtain the quadratic term of Eq. (34) by considering a WKB solution to the wave equation. When following this approach, it is possible to improve the small $\ell$ behavior through the replacement $\ell(\ell+1)\to(\ell+1/2)^{2}$, as shown by Langer Langer (1937), getting the final expression
@@
This is nothing but the standard result for the eikonal scattering between a massless particle [^7] In the eikonal limit, the scalar field scattering amplitude reproduces the gravitational one, while subleading terms encode graviton helicity-flip effects (see comments in the Conclusions). and a massive (lens) field, obtained by resumming all the ladder and cross-ladder diagrams in its Feynman series expansion Abarbanel and Itzykson (1969); Levy and Sucher (1969); ’t Hooft (1987), with the addition of the finite distance Fresnel term. The phase shift in the eikonal limit is then given by Abarbanel and Itzykson (1969); Levy and Sucher (1969); ’t Hooft (1987)
@@
where $\mathcal{M}|_{G}({\bm{q}},s)$ describes the tree-level $t$-channel diagram. In the second line, we have used that $2(s-M^{2})\simeq 4M\omega$, the leading order Lippmann-Schwinger equation (Scattering perspective on gravitational lensing) to write the scattering amplitude in terms of the lens potential, and the Fourier transform conventions in Eq. (19). Thus, comparing to Eq. (9), one derives
@@
reproducing the known result that the two dimensional potential $\psi$ is equal to the gravitational time delay, $\partial_{\omega}\delta$ Schneider et al. (1992). In particular, for the potential of a Newtonian pointlike lens, the eikonal phase shift takes the simple form
@@
which reproduces the matching relation (24) obtained in the Born approximation. This matching clarifies the crucial assumption of Eq. (5), $|\partial_{\tilde{r}}^{2}F|\ll|\omega\partial_{\tilde{r}}F|$, which amounts to adopting the eikonal approximation $\ell\approx\omega b\gg 1$ [together with small deflection angles (4)] as essential in drawing an analogy between the lensing diffraction integral and scattering frameworks, in addition to the condition of classicality listed in Eq. (17).
Finally, we note that the unitarity properties of the scattering amplitude allow one to derive optical-theorem–type statements for the diffraction integral (for work along these lines, but focusing on the consequences of causality, see Tanaka and Suyama (2023); Tanaka et al. (2025); Suyama and Kapadia (2025)). Since, by construction, the eikonal amplitude satisfies the optical theorem, the amplification factor given by the diffraction integral satisfies [^8] Here we are assuming conservative potentials, but this can be easily generalized to include inelastic effects.